# Dynamic Obstacle Avoidance of Unmanned Surface Vehicles in Maritime Environments Using Gaussian Processes Based Motion Planning

### v1

@@ Abstract
During recent years, unmanned surface vehicles are extensively utilised in a variety of maritime applications such as the exploration of unknown areas, autonomous transportation, offshore patrol and others. In such maritime applications, unmanned surface vehicles executing relevant missions that might collide with potential static obstacles such as islands and reefs and dynamic obstacles such as other moving unmanned surface vehicles. To successfully accomplish these missions, motion planning algorithms that can generate smooth and collision-free trajectories to avoid both these static and dynamic obstacles in an efficient manner are essential. In this article, we propose a novel motion planning algorithm named the Dynamic Gaussian process motion planner 2, which successfully extends the application scope of the Gaussian process motion planner 2 into complex and dynamic environments with both static and dynamic obstacles. First, we introduce an approach to generate safe areas for dynamic obstacles using modified multivariate Gaussian distributions. Second, we introduce an approach to integrate real-time status information of dynamic obstacles into the modified multivariate Gaussian distributions. Therefore, the multivariate Gaussian distributions with real-time statuses of dynamic obstacles can be innovatively added into the optimisation process of factor graph to generate an optimised trajectory. The proposed Dynamic Gaussian process motion planner 2 algorithm has been validated in a series of benchmark simulations in the Matrix laboratory and a dynamic obstacle avoidance mission in a high-fidelity maritime environment in the Robotic operating system to demonstrate its functionality and practicability.
@@ 1 Introduction
Benefit from the flexibility and versatility of USVs, they have been widely applied in a series of autonomous maritime missions in the past decades. The employment of USVs in autonomous maritime missions involves three fundamental components:
  • 1.
    First, sensing the surrounding maritime environments through onboard sensors.
  • 2.
    Second, planning safe and smooth trajectories based on sensed information.
  • 3.
    Third, navigating the platforms to follow the generated trajectories to perform relevant missions.
Sensing and navigation are effectively connected through planning, which is responsible for processing sensed information as well as for generating optimised trajectories for navigation. It is important to factor in not just static obstacles such as islands and reefs, but also dynamic obstacles such as other moving USVs during the motion planning processes of USVs. In addition, the motion planning algorithms must plan safe and smooth trajectories within an extreme short time after detecting any dynamic obstacle in maritime environments. Otherwise, USVs may have collided with dynamic obstacles despite that safe and smooth trajectories can be generated afterwards.
To our knowledge, the current mainstream motion planning algorithms are originally designed to generate safe trajectories between start and goal positions while avoiding static obstacles. The fundamental versions of the current mainstream motion planning algorithms do not take the real-time statuses of dynamic obstacles into account when planning trajectories. The current mainstream motion planning algorithms contain: 1) Gird-based algorithms, 2) Sampling-based algorithm, 3) Potential field algorithms, 4) Intelligent algorithms, 5) GP-based algorithms and others. Below is a summary of the current mainstream motion planning algorithms:
  • 1.
    The grid-based algorithms might require a relatively long computation time to find a feasible path due to the nature of the algorithms, especially when the dimension of the motion planning problems increases [1, 2, 3, 4, 5].
  • 2.
    The sampling-based algorithms cannot always guarantee the quality of the generated path. It is possible that the length of the generated path is relatively long. The smoothness of the generated path is also uncertain. Some sampling-based algorithms might require 1) a relatively long time to ensure the convergence because of their randomness and 2) some other time for post-processing path smoothing procedure [1, 2, 6, 7, 8, 9, 10].
  • 3.
    The potential field algorithms are normally limited by the local minima problems. Because of the nature of potential field algorithms, it is impossible to overcome the local minima problem without applying additional strategies [1, 2, 11, 12].
  • 4.
    The intelligent algorithms sometimes require a relatively long computation time and are limited by the local minima problems in some cases [1, 2, 13, 14, 15].
  • 5.
    The GP-based algorithms have several benefits among the mainstream motion planning algorithms [16, 17, 18, 19]. They can generate an optimised path in an efficient manner by smoothing and shortening the path during the planning process. Furthermore, they can eliminate the drawback of local minima by using the probabilistic inference [1, 19].
@@
Almost all the aforementioned mainstream motion planning algorithms have been improved and can avoid dynamic obstacles in some way. A series of examples is provided as follows:
  • 1.
    [20] has proposed an approach to improve the fundamental version of the grid-based motion planning algorithms and enable it to work in complex and dynamic environments.
  • 2.
    [21] has proposed an approach to integrate the real-time status information of dynamic obstacles into a potential field algorithm.
  • 3.
    [22] has proposed an approach to combine a sampling-based motion planning algorithm with the international regulations for preventing collisions at sea to avoid dynamic obstacles.
  • 4.
    [23, 24] have proposed a series of strategies for the sampling-based motion planning algorithms to avoid dynamic obstacles through 1) abandoning the minority of tree branches invalid by dynamic obstacles, 2) maintaining the rest majority of tree branches and 3) generating new collision-free tree branches.
  • 5.
    [25] has proposed a two-layer approach that combines an intelligent algorithm and a potential field algorithm to avoid dynamic obstacles.
  • 6.
    [26] has investigated how to use the GP-based algorithms to avoid dynamic obstacles and proposed a new factor graph structure.
Nevertheless, it still remains to be researched how to use the GP-based motion planning algorithms to overcome static and dynamic obstacles simultaneously. To bridge this research gap, this article has specifically focused on developing a novel motion planning algorithm for USVs to avoid both static and dynamic obstacles in maritime environments. The contributions of this work can be summarised as follows:
  • 1.
    First, an approach is introduced for generating a safe area for a moving USV. More specifically, this is achieved by incorporating the USV dimension specification from relevant sensor data into a multivariate Gaussian distribution’s variance matrix.
  • 2.
    Second, an approach is introduced for integrating various real-time status information of the moving USV into the established safe area. To be more precise, we introduce an approach for integrating the USV speed into the established safe area’s variance matrix. In addition, we design a function to enable both the rotation and translation of the established safe area based upon image processing techniques. Therefore we are able to align it with the course and position of the moving USV.
  • 3.
    Third, a new structure of the used optimisation tool is introduced to deal with the GP-based motion planning problem with both static and dynamic obstacles.
  • 4.
    Fourth, the proposed motion planning algorithm is thoroughly tested and validated in 1) a series of self-constructed simulation environments in MATLAB and 2) a high-fidelity maritime environment in ROS.
@@
In the self-constructed simulation environments in MATLAB, we have tested and validated the performance of the proposed motion planning algorithm in a series of benchmark simulations including:
  • 1.
    A benchmark with different weights of dynamic obstacle likelihood.
  • 2.
    A benchmark with a dynamic obstacle with different velocities.
  • 3.
    A benchmark with dynamic obstacles with different areas.
  • 4.
    A benchmark with multiple dynamic obstacles.
  • 5.
    A benchmark with static and dynamic obstacles.
@@
    The proposed motion planning algorithm can generate a safe and smooth trajectory to avoid dynamic obstacles by considering their areas, directions, speeds and positions.
@@
    The proposed motion planning algorithm can generate an optimised trajectory in an extremely short execution time.
@@
    The proposed motion planning algorithm can be utilised in a variety of complex environments with multiple obstacles and different statuses.
@@
We have also tested and validated the proposed motion planning algorithm in a high-fidelity maritime environment in ROS. In ROS, augmented practicability is achieved based on the consideration of a series of physical properties such as hydrodynamic force, gravity and buoyancy [1, 27]. In comparison with the expensive testing process using a full-sized physical USV, validating the developed motion planning algorithm in ROS can effectively reduce the testing cost and maintain reliable test results simultaneously [1, 2]. A scenario demonstration in the Gazebo simulation environment in ROS is provided in Fig. 1.
The reminder of this article is organised as follows. Section 2 presents the mathematical model of the GPMP2 algorithm in detail. Section 3 presents the proposed D-GPMP2 algorithm in detail. Section 4 presents a new structure of the optimisation tool used in the research. Section 5 demonstrates a series of simulation results of the proposed motion planning algorithm in MATLAB. Section 6 demonstrates the practical performance of the proposed motion planning algorithm in ROS, while Section 7 concludes this article and proposes some future research directions.
@@ 2 GPMP2
This section presents the traditional GP-bsaed motion planning algorithm (GPMP2 algorithm) and summarises the previous work in [1, 2, 16, 17, 18, 19]. In general, the GPMP2 algorithm considers a motion planning problem as a standard optimisation problem. To be more specific, the GPMP2 algorithm views the potential sample trajectories as a GP prior, a conditional constraint that facilitates the capability of avoiding static obstacles as a static obstacle likelihood and the optimisation process as a MAP estimation to obtain an optimal solution [1, 19].
@@ 3 D-GPMP2
This section presents the proposed D-GPMP2 algorithm, in detail, which includes two new parts added into the original GPMP2 algorithm: 1) Dynamic obstacle area modelling part and 2) Dynamic obstacle velocity modelling part.
@@
For the dynamic obstacle area modelling, we propose a method that combines the AIS data and multivariate Gaussian distribution to construct a safe area for the target USV.
@@ 3.1 Dynamic obstacle area modelling
The AIS is a fundamental equipment mounted on the USVs to display maritime traffic information and communicate with other USVs. In general, maritime traffic information information from the AIS contains: 1) the target USV position from the GPS, 2) the target USV dimension, 3) the target USV speed, 4) the target USV course (also known as target USV direction) and others [32].
where $\zeta$ is an adjustable positive scaling factor, $[A+B]$ represents the length of the USV, $[C+D]$ represents the width of the USV and the USV safe area is constructed by a multivariate Gaussian distribution. Furthermore, the self-defined safe radius $R$ should ensure the multivariate Gaussian distribution at least fully covers the USV.
Figure 4: (a) The safe area of the Heron USV in MATLAB from the 2-D perspective. (b) The safe area of the Heron USV in MATLAB from the 3-D perspective. (c) The safe area of the WAM-V 20 USV in MATLAB from the 2-D perspective. (d) The safe area of the WAM-V 20 USV in MATLAB from the 3-D perspective. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
@@ 3.2 Dynamic obstacle velocity modelling
$M_{N\times N}=f_{crs\_pos}(M_{N\times N},\alpha,x_{gps},y_{gps}),$ (21)
@@
Fig. 6 provides a series of examples to give a more intuitive understanding of the working process of $f_{crs\_pos}$, while Fig. 7 demonstrates the modelling results of target USVs with different velocities and positions based on (20) and (21).
@@ 3.5 MAP estimation
Figure 9: Overall structure of the factor graph. $f^{p}_{0}(\theta_{0})$ and $f^{p}_{N}(\theta_{N})$ represent the regular prior factors, $f^{gp}_{i}(\theta_{i},\theta_{i+1})$ represent the GP prior factors, $f^{st\_obs}_{i}(\theta_{i})$ represent the static obstacle likelihood factors, $f^{st\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$ represent the interpolated static obstacle factors, $f^{dy\_obs}_{i}(\theta_{i})$ represent the dynamic obstacle likelihood factors and $f^{dy\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$ represent the interpolated dynamic obstacle factors.

### v2

@@ Abstract
During recent years, unmanned surface vehicles are extensively utilised in a variety of maritime applications such as the exploration of unknown areas, autonomous transportation, offshore patrol and others. In such maritime applications, unmanned surface vehicles executing relevant missions that might collide with potential static obstacles such as islands and reefs and dynamic obstacles such as other moving unmanned surface vehicles. To successfully accomplish these missions, motion planning algorithms that can generate smooth and collision-free trajectories to avoid both these static and dynamic obstacles in an efficient manner are essential. In this article, we propose a novel motion planning algorithm named the Dynamic Gaussian process motion planner 2, which successfully extends the application scope of the Gaussian process motion planner 2 into complex and dynamic environments with both static and dynamic obstacles. First, we introduce an approach to generate safe areas for dynamic obstacles using modified multivariate Gaussian distributions. Second, we introduce an approach to integrate real-time status information of dynamic obstacles into the modified multivariate Gaussian distributions. The multivariate Gaussian distributions with real-time statuses of dynamic obstacles can be innovatively added into the optimisation process of factor graph to generate an optimised trajectory. We also develop a variant of the proposed algorithm that integrates the international regulations for preventing collisions at sea, enhancing its operational effectiveness in maritime environments. The proposed algorithms have been validated in a series of benchmark simulations and a dynamic obstacle avoidance mission in a high-fidelity maritime environment in the Robotic operating system to demonstrate the functionality and practicability.
@@ 1 Introduction
COLREG Convention on the international regulations for preventing collisions at sea
@@
D-GPMP2-COLREGs Dynamic Gaussian process motion planner 2 based on COLREG rules
DWA Dynamic window approach
@@
Owing to the inherent flexibility and operational versatility, USVs have been extensively deployed in diverse autonomous maritime missions over recent decades. The deployment of USVs for autonomous maritime operations comprises three essential elements: first, sensing the surrounding maritime environments through onboard sensors, second, planning safe and smooth trajectories based on sensed information and third, navigating the platforms to follow the generated trajectories to perform relevant missions. Sensing and navigation are effectively connected through planning, which is responsible for processing sensed information as well as for generating optimised trajectories for navigation. It is important to factor in not just static obstacles such as islands and reefs, but also dynamic obstacles such as other moving USVs during the motion planning processes of USVs. Furthermore, motion planning algorithms are required to plan safe and smooth trajectories within an extreme short time such as 100 [ms] after detecting any dynamic obstacle in maritime environments. Otherwise, the USV may already have collided with dynamic obstacles during the planning latency period, even if subsequently generated trajectories are both safe and smooth.
Current mainstream motion planning algorithms contain: 1) gird-based algorithms [1, 2, 3, 4, 5], 2) sampling-based algorithm [1, 2, 6, 7, 8, 9, 10], 3) potential field algorithms [1, 2, 11, 12], 4) intelligent algorithms [1, 2, 13, 14, 15], 5) GP-based algorithms [16, 17, 18, 19] and others. They can also be divided into two categories: 1) path planning methods and 2) trajectory planning methods. In path planning, a geometric path is defined as a sequence of waypoints focusing solely on spatial coordinates. In trajectory planning, the geometric path is further augmented with time-dependent information. Except traditional mainstream motion planning algorithms, obstacle avoidance strategy can dynamically adjust the path or trajectory and directly generates control commands based on real-time sensor data during the USV’s movement to prevent collisions with obstacles. The GP-based algorithms have several benefits among the mainstream motion planning algorithms [16, 17, 18, 19]. They can generate an optimised path at an extremely fast speed while smoothing and shortening the path during the planning process. Furthermore, they can eliminate the drawback of local minima by using the probabilistic inference [1, 19]. The benefits of GP-based algorithms compared with mainstream motion planning algorithms and obstacle avoidance strategies are summarised in Table 2.
Table 2: Benefits of GP-based algorithms compared with mainstream motion planning algorithms and obstacle avoidance strategies.
Methods Benefits
GP-based algorithms 1. Short execution time 2. Smooth path 3. No local minima issue
To ensure accuracy, the effects of both static obstacles and dynamic obstacles should be modelled in the motion planning problems of USVs. Nearly all the mainstream motion planning algorithms have been improved and can avoid dynamic obstacles. There are many studies about using grid-based motion planning algorithms to avoid dynamic obstacles [20, 21, 22, 23]. For example, [20] has proposed an approach to improve the fundamental version of the grid-based motion planning algorithms and enable it to work in complex and dynamic environments. [21] proposes an improved D* Lite algorithm, a grid-based motion planning algorithm designed for marine robots such as USVs to navigate dynamic maritime environments while avoiding moving obstacles. Similar to grid-based motion planning algorithms, there are also many studies about using sampling-based motion planning algorithms to avoid dynamic obstacles [24, 25, 26, 27, 28]. For example, [24, 25] have proposed different approaches to combine sampling-based motion planning algorithm with COLREG collision avoidance regulations. Furthermore, [26, 27] have proposed a series of strategies for the sampling-based motion planning algorithms to avoid dynamic obstacles through 1) abandoning the minority of tree branches invalid by dynamic obstacles, 2) maintaining the rest majority of tree branches and 3) generating new collision-free tree branches. Potential field motion planning algorithms and intelligent motion planning algorithms have been researched for avoiding dynamic obstacles for many years as well [29, 30, 31, 32]. For example, [29] has proposed an approach to integrate the real-time status information of dynamic obstacles into a potential field algorithm. [30] has proposed a two-layer approach that combines an intelligent algorithm and a potential field algorithm to avoid dynamic obstacles. [33] employs an obstacle avoidance strategy (DWA) as a local motion planner to generate collision-free paths for USVs navigating among dynamic obstacles.
Despite some prior studies has modified some mainstream motion planning algorithms and obstacle avoidance strategies to handle dynamic obstacles, there was not many studies applying the state-of-the-art trajectory optimisation method (GP-based algorithms) to deal with the motion planning problems of avoiding static and dynamic obstacles simultaneously. Although [34] has investigated how to use the GP-based algorithms to avoid dynamic obstacles by proposing a new factor graph structure. But this is the only published research about this topic and it just adjusted the factor graph structure without providing a specific method to model the dynamic obstacles. To bridge this research gap, this article has specifically focused on developing a novel motion planning algorithm for USVs to avoid both static and dynamic obstacles in maritime environments.
@@
In the self-constructed simulation environments in MATLAB, we have tested and validated the performance of the proposed motion planning algorithms in a series of qualitative and quantitative benchmark simulations including 1) a benchmark with different weights of dynamic obstacle likelihood, 2) a benchmark with a dynamic obstacle with different velocities, 3) a benchmark with dynamic obstacles with different areas, 4) a benchmark with multiple dynamic obstacles, 5) a benchmark with static and dynamic obstacles, 6) a benchmark with a current existing obstacle avoidance method for quantitative comparison and 7) a COLREGs-compliant benchmark simulation.
@@
    The proposed D-GPMP2 algorithm can generate a safe and smooth trajectory to avoid dynamic obstacles by considering their areas, directions, speeds and positions.
@@
    The proposed D-GPMP2 algorithm can generate an optimised trajectory in an extremely short execution time.
@@
    The proposed D-GPMP2 algorithm can be utilised in a variety of complex environments with multiple obstacles and different statuses.
  • 4.
    The proposed D-GPMP2-COLREGs algorithm effectively performs dynamic obstacle avoidance in maritime environments while maintaining compliance with COLREGs regulations.
@@
We have also tested and validated the proposed D-GPMP2 algorithm in a high-fidelity maritime environment in ROS. In ROS, augmented practicability is achieved based on the consideration of a series of physical properties such as hydrodynamic force, gravity and buoyancy [1, 35]. In comparison with the expensive testing process using a full-sized physical USV, validating the developed motion planning algorithm in ROS can effectively reduce the testing cost and maintain reliable test results simultaneously [1, 2]. A scenario demonstration in the Gazebo simulation environment in ROS is provided in Fig. 1.
The reminder of this article is organised as follows. Section 2 presents the mathematical model of the GPMP2 algorithm in detail. Section 3 presents the proposed D-GPMP2 algorithm in detail. Section 4 presents the proposed D-GPMP2-COLREGs algorithm in detail. Section 5 presents a new structure of the optimisation tool used in the research. Section 6 demonstrates a series of simulation results of the proposed motion planning algorithms in MATLAB. Section 7 demonstrates the practical performance of the proposed D-GPMP2 algorithm in ROS, while Section 8 concludes this article and proposes some future research directions.
@@ 2 GPMP2
This section presents the traditional GP-bsaed motion planning algorithm (GPMP2 algorithm) and summarises the previous work in [1, 2, 16, 17, 18, 19]. In general, the GPMP2 algorithm considers a motion planning problem as a standard optimisation problem. To be more specific, the GPMP2 algorithm views the potential sample trajectories as a GP prior, a conditional constraint that facilitates the capability of avoiding static obstacles as a static obstacle likelihood and the optimisation process as a MAP estimation to obtain an optimal solution [1, 19]. The trajectories generated by GPMP2 are 1) relatively short (as GPMP2 incrementally optimise the posterior during the motion planning process) and 2) smooth (as the trajectories are time-continuous and differentiable).
@@ 3 D-GPMP2
This section presents the proposed D-GPMP2 algorithm, in detail, which includes two new parts added into the original GPMP2 algorithm: 1) dynamic obstacle area modelling part and 2) dynamic obstacle velocity modelling part. Compared with previous studies of using multivariate Gaussian distributions to model dynamic obstacles, the novelties of our method are: 1) a specific way is proposed to integrate sensor data such as AIS data, camera data and other data into the dynamic obstacle modelling and 2) our method consider both the area and velocity of a dynamic obstacle during the modelling process.
@@
For the dynamic obstacle area modelling, we propose a method that combines the sensor data such as AIS data, camera data and other data and multivariate Gaussian distribution to construct a safe area for the target USV.
@@ 3.1 Dynamic obstacle area modelling
Dynamic obstacle area modelling method used in this article automatically adjusts model boundaries to reflect real-world obstacle sizes. Different sensors can be installed on a USV to collect information about nearby obstacles. For example, the AIS is a fundamental equipment mounted on the USVs to display maritime traffic information and communicate with other USVs. According to the official website of the USV platform used in this article, the WAM-V USV platform offers factory-configured sensors, such as AIS and other options [40]. In general, maritime traffic information information from the AIS contains: 1) the target USV position from the GPS, 2) the target USV dimension, 3) the target USV speed, 4) the target USV course (also known as target USV direction) and others [41]. When the AIS is not functioning properly, alternative sensors such as cameras, event cameras and GPS can be utilised to obtain similar information.
Figure 4: (a) Heron USV in virtual environment [42]. (b) The safe area of the Heron USV from the 2-D perspective. (c) The safe area of the Heron USV from the 3-D perspective. (c) WAM-V 20 USV in virtual environment [35]. (d) The safe area of the WAM-V 20 USV from the 2-D perspective. (e) The safe area of the WAM-V 20 USV from the 3-D perspective. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
@@
where $\zeta$ is an adjustable positive scaling factor and it is used for experimental calibration, $[A+B]$ represents the length of the USV, $[C+D]$ represents the width of the USV and the USV safe area is constructed by a multivariate Gaussian distribution. Furthermore, the self-defined safe radius $R$ should ensure the multivariate Gaussian distribution at least fully covers the USV.
@@ 3.2 Dynamic obstacle velocity modelling
$M^{\prime}_{N\times N}=f_{crs\_pos}(M_{N\times N},\alpha,x_{gps},y_{gps}),$ (21)
@@
Fig. 6 provides a series of examples to give a more intuitive understanding of the working process of $f_{crs\_pos}$, while Fig. 7 demonstrates the modelling results of target USVs with different velocities and positions based on (20) and (21). The detailed process of constructing dynamic obstacle model is provided in Algorithm 1.
1 Input: Length of the target USV ($l$), width of the target USV ($w$), speed of the target USV ($v$), position of the target USV [$x_{gps}$, $y_{gps}$] and the difference between the target USV course and the positive half axis of $Y$ axis ($\alpha$)
2 Compute the safe radius of the target USV by using equation (12);
3 Perform dynamic area modelling for the target USV by using equation (16) and (17);
4 Perform dynamic velocity modelling for the target USV by using equation (19);
5 Calibrate the constructed dynamic velocity model by using equation (20);
6 Adjust the position and orientation of the calibrated dynamic obstacle model by using equation (21);
7 Output: Dynamic obstacle likelihood l($\theta$; $e_{dy\_obs}$)
Algorithm 1 Dynamic obstacle modelling
@@ 3.5 MAP estimation
4 D-GPMP2-COLREGs
This section extends the proposed D-GPMP2 algorithm through COLREG rules integration and proposes a variant named D-GPMP2-COLREGs algorithm.
The COLREG rules define behaviours for USVs in overtaking, head-on and crossing scenarios [46]:
  • 1.
    Overtaking: the overtaking USV is required to keep out of the way of the USV that is being overtaken.
  • 2.
    Head-on: if two USVs are meeting on reciprocal or nearly reciprocal courses, each USV is required to adjust its course to starboard and passes on the port side of the other.
  • 3.
    Crossing: if two USVs are crossing, the USV that has the other on its own starboard side is required to keep out of the way. This USV is also required to avoid crossing ahead of the other USV.
D-GPMP2-COLREGs algorithm is proposed to enable the proposed D-GPMP2 algorithm to generate trajectories that comply with COLREGs rules. The pseudocode of D-GPMP2-COLREGs is detailed in Algorithm 2: colregs_region() is a function that calculates feasible regions for trajectory planning, ensuring compliance with COLREGs rules. random_angle() is a function that generates a uniformly distributed random angle within given bounds. factor_graph() constructs a factor graph based on a series of acquired data. levenberg_marquardt() is a function that performs gradient-based optimisation via the Levenberg Marquardt method. The feasible regions for D-GPMP2-COLREGs algorithm to generate trajectories in overtaking, head-on, and crossing scenarios in colregs_region() are detailed below:
1 Input: Start state $\theta_{0}$, goal state $\theta_{N}$ and dynamic obstacle position [$x_{gps}$, $y_{gps}$]
2 Precompute static obstacle likelihood l($\theta$; $e_{st\_obs}$) and dynamic obstacle likelihood l($\theta$; $e_{dy\_obs}$)
3 S = colregs_region($\theta_{0}$, [$x_{gps}$, $y_{gps}$]);
4 $\psi$ = random_angle($\theta_{0}$, S);
5 $G$ = factor_graph($\theta_{0}$, $\theta_{N}$, $\psi$, l($\theta$; $e_{st\_obs}$), l($\theta$; $e_{dy\_obs}$));
6 $\theta^{*}$ = levenberg_marquardt($G$);
7 Output: Optimal path $\theta^{*}$
Algorithm 2 D-GPMP2-COLREGs (Dynamic Gaussian process motion planner 2 based on COLREGs rules)
Figure 9: The regions that D-GPMP2-COLREGs algorithm generates trajectories in overtaking, head-on and crossing scenarios. (a) overtaking scenario, (b) head-on scenario and (c) crossing scenario. The blue triangle indicates our USV, while the orange triangle represents the encountered USV. The shaded regions represents the regions that D-GPMP2-COLREGs algorithm can generate trajectories.
  • 1.
    Overtaking scenario: trajectories can be generated on either sides of the encountered USV as demonstrated in Fig. 9(a).
  • 2.
    Head-on scenario: trajectories can be generated on the starboard side of our USV as demonstrated in Fig. 9(b).
  • 3.
    Crossing scenario: trajectories can be generated on the starboard side of our USV and the port side of the encountered USV as demonstrated in Fig. 9(c).
Since the coordinate systems in the real world and in the high-fidelity virtual world in Gazebo differ, it is required to calibrate the error between these coordinate systems before calculating any angles, as we have done in [2].
Figure 10: Overall structure of the factor graph. $f^{p}_{0}(\theta_{0})$ and $f^{p}_{N}(\theta_{N})$ represent the regular prior factors, $f^{gp}_{i}(\theta_{i},\theta_{i+1})$ represent the GP prior factors, $f^{st\_obs}_{i}(\theta_{i})$ represent the static obstacle likelihood factors, $f^{st\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$ represent the interpolated static obstacle factors, $f^{dy\_obs}_{i}(\theta_{i})$ represent the dynamic obstacle likelihood factors and $f^{dy\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$ represent the interpolated dynamic obstacle factors. $f^{st\_obs}_{i}(\theta_{i})$ and $f^{st\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$ contains static obstacle information from the surrounding environment. $f^{dy\_obs}_{i}(\theta_{i})$ and $f^{dy\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$ contains dynamic obstacle information from the surrounding environment.